I don't know how to solve this and I can't find anything in my text for this...help please =/
The weight in kilograms,W, of a female African elephant at age t years is given by W(t)=2600(1-0.51e^-0.075t). Estimate the age, to the nearest year, of an adult female African elephant weighing 1800 kilograms.

Answer by Cromlix(4381)

W(t) = 2600(1 - 0.51e^-0.075t)
1800 = 2600(1 - 0.51e^-0.075t
1800/2600 = 1 - 0.51e^-0.075t
1800/2600 - 1 = -0.51e^-0.075t
(1800/2600 - 1)/-0.51 = e^-0.075t
0.6033 = e^-0.075t
ln 0.6033 = -0.075t ln e
(ln e = 1)
ln 0.6033/-0.075 = t
t = 6.7 years
Hope this helps.
:-)